OpenStudy (anonymous):

need help finding this derivative without the quotient rule. -7t^3 + 3/(t^3) - 3/(3rdrt(t))

OpenStudy (anonymous):

is this the function: \[-7t^3 + \frac{3}{t^3} -\frac{3}{\sqrt[3]{t}}\]

OpenStudy (anonymous):

or this : \[\LARGE -7t^3+\frac{3}{t^3}-\frac{3}{3\sqrt t}\] ?

OpenStudy (shayaan_mustafa):

then use difference quotient or you can say that first principle.

OpenStudy (shayaan_mustafa):

if quotient rule is not the choice then use above specified rule. good luck.

OpenStudy (anonymous):

yea eig that is the correct equation

OpenStudy (anonymous):

\[f(x) = -7t^3 +\frac{3}{t^3} -\frac{3}{\sqrt[3]{t}} = -7t^3 +3t^{-3} -3t^{-\frac{1}{3}} \]

OpenStudy (anonymous):

differentiate the normal way ax^n goes to nax^(n-1)

OpenStudy (anonymous):

the only confusion i had was converting fractions into exponents. if its 3/ (t^3 that just becomes 3t^-3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and obviously the other is 1/3 because of the root. Thanks very much that was my only confusion

OpenStudy (anonymous):

no problem :) happy mathsing